# Computational Bifurcation Analysis

### v1

@@ 2. Illustration: Continuous Stirred Tank Reactor
Equilibria of (1) are pairs of values of $x$ and $y$, given allowable values for the parameters $\beta$, $\delta$, $\gamma$, and $\sigma$, for which $\dot{x}$ and $\dot{y}$ both equal 00. By elimination, a pair $(x,y)$ is an equilibrium provided that $x=\delta y/\sigma$, where
@@
i.e., by necessity that $0<y<\sigma/\delta$ and $10<x<$ 1, since the right-hand side is non-negative. From an analysis of the graph of the right-hand side of (2) as a function of $y$ (say an exploration of local maxima and minima and limiting behaviors as $y\rightarrow 0$ and $y\rightarrow\infty$) and recognizing the graph of the left-hand side as a straight line with positive ordinate intercept and negative slope, it follows that there exists at least one equilibrium for any allowable choice of $\beta$, $\delta$, $\gamma$, and $\sigma$. The equilibrium is unique provided that either $\beta\geq 1/4$, $\delta>(1-4\beta)/e^{2}$, or
@@ 2.1. The case when $\beta=0$
[…]
@@
which add to 00 along a curve with
@@
(where $\operatorname{trace}J=0$) and one of which equals 00 along a curve with
@@
with $p(x)$ given above. It follows that $J^{2}$ is a multiple of the $2\times 2$ identity matrix along the curve (5), specifically that $J^{2}=-kI_{2}$ along this curve. As expected, $k>0$ for $x\in(x_{\mathrm{BT},1},x_{\mathrm{BT},2})$, such that those solutions correspond to Hopf bifurcations. On the other hand, values of $x$ with $k<0$ correspond to neutral saddles (for which the two eigenvalues of $J$ are both real and sum to 00).
@@
along which $k$ is always non-positive. Along this family, $-\sqrt{-k}$ is an eigenvalue of $J$ with eigenvector $(\cos\theta,\sin\theta)^{\mathsf{T}}$ while the other eigenvalue equals $e^{y}(x\cot\theta-\gamma)/x\gamma$. For a given $\theta$, the two families in (7) and (8) intersect on a neutral saddle with
@@
With reference to Fig. 2, by solving the equation $\ell_{1}=0$ for $\gamma$ while recalling that $\gamma<x^{2}(1-x)$ (since $p(x)<0$), one concludes that the first Lyapunov coefficient vanishes for $x_{\mathrm{BT},1}<x<x_{\mathrm{BT},2}$ at a unique value of $x$ when $21/8\leq\gamma<(7-3\sqrt{5})/$ 2, at two distinct values of $x$ when $80<\gamma<1/$ 8, and for no value of $x$ when $\gamma\geq(7-3\sqrt{5})/2$. Such loci of vanishing first Lyapunov coefficient, denoted by the subscript DH, correspond to degenerate Hopf bifurcations, such that the Hopf bifurcations on either side are supercritical where $\ell_{1}<0$ and subcritical where $\ell_{1}>0$. As $\gamma$ approaches $(7-3\sqrt{5})/2$ from below, the single $x_{\mathrm{DH}}\rightarrow x_{\mathrm{BT},1}$. Similarly, the left-most $x_{\mathrm{DH}}\rightarrow x_{\mathrm{BT},1}$ as $\gamma$ approaches $1/8$ from below. Away from these degenerate values of $\gamma$, it is evident from the closed-form expression that $\left|\ell_{1}\right|$ grows without bound as $x$ approaches $x_{\mathrm{BT},1}$ and $x_{\mathrm{BT},2}$.
@@ 2.2. Numerical bifurcation analysis
$[0,1)\ni s\mapsto\left(\frac{se^{-\frac{s(1-s)}{s(1-s)-\gamma}}}{s(1-s)-\gamma},\frac{se^{-\frac{s(1-s)}{s(1-s)-\gamma}}}{1-s}\right)$
@@
The continuation algorithm terminates with a failure to converge (denoted by MX) as, near $\gamma=1/8$, $\ell_{1}$ varies rapidly between 00 at the DH point and $\infty$ at the BTP point (which are coincident in the limit of $\gamma=1/8$).
@@ 2.3. Periodic orbits
Figure 6. Bifurcation diagram for equilibria and periodic orbits with free parameters $\sigma$ and $\delta$ in a small parameter region containing the Bogdanov-Takens point $(\sigma_{\mathrm{BT,1}},\delta_{\mathrm{BT,1}})$ (subscript $1$ not shown in the figure) in Fig. 3(b). For better visibility of the bifurcation curves, the coordinates of the parameter plane have been rotated counterclockwise by $\theta=150^{\circ}$ around the Bogdanov-Takens point, such that $\tilde{\sigma}=\cos\theta(\sigma-\sigma_{\mathrm{BT}})-\sin\theta(\delta-\delta_{\mathrm{BT}})$, $\tilde{\delta}=\sin\theta(\sigma-\sigma_{\mathrm{BT}})+\cos\theta(\delta-\delta_{\mathrm{BT}})$. Other parameters: $\beta=0$, $\gamma=0.1$.
For homoclinic orbits connecting to saddle-node bifurcation equilibria, the determinant (det) should equal 00, whereas the determinant is negative in the saddle case. Across the transition between such orbits, the shape of the homoclinic turns from a smooth curve to a non-smooth curve in phase space (the “corner” is the location of the saddle, see phase portraits in Fig. 5). At the transition point, the homoclinic orbit is a non-central connection to a saddle-node [2]. In the screen output, approximate detection of such a transition is accomplished by monitoring for a crossing by det of $-0.05$ (a suitably chosen small, negative number) and the corresponding points are denoted by NCS. Four such non-central connections to saddle-nodes are found near the Bogdanov-Takens point $(\sigma_{\mathrm{BT,1}},\delta_{\mathrm{BT,1}})$ at the end points of the segments labeled SNIC in Fig. 6. Magnification (not included here but available using the provided code) shows that the homoclinic curve is smooth across these points.
@@ 3.1. Algebraic problems
[…]
where $J=\partial_{x}f(x,p)$. (Note that the numerical implementation does not perform such elimination since if $J$ has large entries, entries in $J^{2}$ may be even larger.) Suppose that a vector $\hat{v}_{1}$ satisfies the first equation in (9). Then $-k$ must be an eigenvalue of $J^{2}$ with eigenvector $\hat{v}_{1}$. If $k>0$, then $\hat{v}_{1}$ cannot be an eigenvector also of $J$ (since $-k<0$). It follows that $\hat{v}_{1}$ and $J\hat{v}_{1}$ span a two-dimensional invariant subspace of $J$ for the purely imaginary eigenvalue $\sqrt{-k}$ with corresponding complex eigenvector $\sqrt{-k}\hat{v}_{1}+J\hat{v}_{1}$. In particular, there exist locally unique scalars $\alpha$ and $\beta$ such that the vector $\alpha\hat{v}_{1}+\beta J\hat{v}_{1}$ satisfies all three of the equations in (9).
Suppose, instead, that $k<0$. Then, if $\hat{v}_{1}$ is not an eigenvector of $J$, $\hat{v}_{1}$ and $J\hat{v}_{1}$ span an invariant subspace of $J^{2}$, which is also an invariant subspace for $J$. It follows that $J$ must have the real eigenvalues $\sqrt{-k}$ and $-\sqrt{-k}$, i.e., the corresponding equilibrium is a neutral saddle. There again exist locally unique scalars $\alpha$ and $\beta$ such that the vector $\alpha\hat{v}_{1}+\beta J\hat{v}_{1}$ satisfies all three of the equations in (9).
If, instead, $\hat{v}_{1}$ is an eigenvector of $J$ (by necessity corresponding to an eigenvalue $\sqrt{-k}$ or $-\sqrt{-k}$), then $\hat{v}_{1}$ and $J\hat{v}_{1}$ do not span a two-dimensional invariant subspace for $J$ and there are no constraints on the other eigenvalues of $J$. Orthogonality with respect to $\xi$ here plays the role of the eigenvalue conditions in the preceding two paragraphs. In the CSTR model example, this case corresponds to the solution family (8). Such a solution family may intersect a family of neutral saddles at a unique branch point, as observed in the preceding section for the CSTR model example.
@@ 3.2. Boundary-value problems
The function $\tau\mapsto x(\tau)$ is next approximated by a continuous function $\tau\mapsto x_{\mathrm{d}}(\tau)$ with $L$ polynomial pieces on subintervals defined on a mesh $10=\tau_{0}<\tau_{1}<\ldots\tau_{L}=$ 1, each of degree $n_{\mathrm{deg}}$. Finally, let $\tau_{\mathrm{coll},j}$ for $j=1,\ldots,n_{\mathrm{deg}}$ be collocation parameters defining the discretization method (e.g., Gauss-Legendre nodes [12] scaled to a base interval $[0,1]$) and $P_{\mathrm{d}}$ be the projection that maps the function $\tau\mapsto y(\tau)$ onto the unique (possibly discontinuous) piecewise-polynomial function $\tau\mapsto[P_{\mathrm{d}}y](\tau)$ that is of degree $n_{\mathrm{deg}}-1$ on each mesh interval and satisfies
@@
The defining problem (12) for a saddle-node bifurcation of periodic orbits is guided by the desire to ensure that the monodromy matrix associated with the variational problem has a generalized eigenvector $y_{\mathrm{d}}(0)$ corresponding to the Floquet multiplier $1$, orthogonal to $\dot{x}_{\mathrm{d}}(0)=f(x_{\mathrm{d}}(0),p)$. In a similar fashion, other generic local bifurcations of periodic orbits, such as period-doubling bifurcations or torus bifurcations have defining problems that ensure that the monodromy matrix has an eigenvector for a desired Floquet multiplier. For period-doubling bifurcations, this is achieved by omitting the fifth component in (12) and replacing the next-to-last component with $y_{\mathrm{d}}(1)+v$, such that $y_{\mathrm{d}}(0)$ is an eigenvector corresponding to a Floquet multiplier $-1$.
@@ 3.3. Periodic orbits near homoclinics
Consider, first, the scenario shown in Fig. 5, in which a family of periodic orbits accumulate on a connecting orbit to a saddle-node bifurcation equilibrium at a saddle-node-on-invariant curve bifurcation as $\sigma\downarrow\sigma_{\mathrm{snic}}$. Here, increases in the period of nearby periodic orbits are dominated by a slowing down of the dynamics in the immediate vicinity of the ghost of the saddle-node bifurcation equilibrium as represented schematically in Fig. 9(a). This depicts a size-$\varepsilon$ neighborhood of the saddle-node bifurcation equilibrium in the center manifold (horizontal axis is $\sigma$, vertical axis is the other center manifold direction, called $x_{\mathrm{c}}$). The parabolic curve in Fig. 9(a) is the family of equilibria associated with the saddle-node bifurcation (red dashed unstable, and green solid stable). Expanding the vector field to second order about the saddle-node bifurcation, one concludes that the time for $x_{\mathrm{c}}$ to increase from $-\varepsilon$ to $\varepsilon$ for $\sigma>\sigma_{\mathrm{snic}}$ varies as
@@ 4.1. A Network of Chemical Oscillators
For fixed $A$ and $\delta$, such that $11A^{2}(11-9\delta)>100\delta^{2}$, and under variations in $\epsilon$, the second pair of eigenvalues, $\kappa_{2,\pm}$, crosses the imaginary axis transversally for $\epsilon=\epsilon^{*}:=\delta/44$ at ±$\pm\mathrm{i}\omega$ with
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
... 1 EP 1.4473e+01 5.9863e+00 1.8536e+00
... 2 BP 1.9650e+01 7.2452e+00 2.6173e+00
... 3 UST 1.9650e+01 7.2452e+00 2.6173e+00
... 4 BP 2.2810e+01 8.5430e+00 3.6744e+00
... 5 UST 2.2810e+01 8.5430e+00 3.6743e+00
... 6 UST 2.5346e+01 1.0683e+01 4.8199e+00
... 7 FP 2.5346e+01 1.0683e+01 4.8199e+00
... 8 BP 1.5355e+01 1.2014e+01 4.0979e+00
... 9 UST 1.5355e+01 1.2014e+01 4.0979e+00
... 10 UST 1.5209e+01 1.2009e+01 4.0666e+00
... 11 BP 1.5209e+01 1.2009e+01 4.0666e+00
... 12 BP 1.4616e+01 1.1820e+01 3.9017e+00
... 13 UST 1.4616e+01 1.1820e+01 3.9017e+00
... 14 EP 1.4000e+01 1.1364e+01 3.6991e+00
[…]

### v2

@@ 2. Illustration: Continuous Stirred Tank Reactor
Equilibria of (1) are pairs of values of $x$ and $y$, given allowable values for the parameters $\beta$, $\delta$, $\gamma$, and $\sigma$, for which $\dot{x}$ and $\dot{y}$ both equal 0. By elimination, a pair $(x,y)$ is an equilibrium provided that $x=\delta y/\sigma$, where
@@
i.e., by necessity that $0<y<\sigma/\delta$ and $0<x<1$, since the right-hand side is non-negative. From an analysis of the graph of the right-hand side of (2) as a function of $y$ (say an exploration of local maxima and minima and limiting behaviors as $y\rightarrow 0$ and $y\rightarrow\infty$) and recognizing the graph of the left-hand side as a straight line with positive ordinate intercept and negative slope, it follows that there exists at least one equilibrium for any allowable choice of $\beta$, $\delta$, $\gamma$, and $\sigma$. The equilibrium is unique provided that either $\beta\geq 1/4$, $\delta>(1-4\beta)/e^{2}$, or
@@ 2.1. The case when $\beta=0$
(3) $\left\{\left(x,y,\delta,\sigma\right)\,|\;0<x<1,\delta=\frac{xe^{-y}}{1-x},\,\sigma=\frac{ye^{-y}}{1-x}\right\}.$
@@
which add to 0 along a curve with
@@
(where $\operatorname{trace}J=0$) and one of which equals 0 along a curve with
@@
with $p(x)$ given above. It follows that $J^{2}$ is a multiple of the $2\times 2$ identity matrix along the curve (5), specifically that $J^{2}=-kI_{2}$ along this curve. As expected, $k>0$ for $x\in(x_{\mathrm{BT},1},x_{\mathrm{BT},2})$, such that those solutions correspond to Hopf bifurcations. On the other hand, values of $x$ with $k<0$ correspond to neutral saddles (for which the two eigenvalues of $J$ are both real and sum to 0).
@@
along which $k$ is always non-positive. Along this family, $\sqrt{-k}$ is an eigenvalue of $J$ with eigenvector $(\cos\theta,\sin\theta)^{\mathsf{T}}$ while the other eigenvalue equals $e^{y}(x\cot\theta-\gamma)/x\gamma$. For a given $\theta$, the two families in (7) and (8) intersect on a neutral saddle with
@@
With reference to Fig. 2, by solving the equation $\ell_{1}=0$ for $\gamma$ while recalling that $\gamma<x^{2}(1-x)$ (since $p(x)<0$), one concludes that the first Lyapunov coefficient vanishes for $x_{\mathrm{BT},1}<x<x_{\mathrm{BT},2}$ at a unique value of $x$ when $1/8\leq\gamma<(7-3\sqrt{5})/2$, at two distinct values of $x$ when $0<\gamma<1/8$, and for no value of $x$ when $\gamma\geq(7-3\sqrt{5})/2$. Such loci of vanishing first Lyapunov coefficient, denoted by the subscript DH, correspond to degenerate Hopf bifurcations, such that the Hopf bifurcations on either side are supercritical where $\ell_{1}<0$ and subcritical where $\ell_{1}>0$. As $\gamma$ approaches $(7-3\sqrt{5})/2$ from below, the single $x_{\mathrm{DH}}\rightarrow x_{\mathrm{BT},1}$. Similarly, the left-most $x_{\mathrm{DH}}\rightarrow x_{\mathrm{BT},1}$ as $\gamma$ approaches $1/8$ from below. Away from these degenerate values of $\gamma$, it is evident from the closed-form expression that $\left|\ell_{1}\right|$ grows without bound as $x$ approaches $x_{\mathrm{BT},1}$ and $x_{\mathrm{BT},2}$.
@@ 2.2. Numerical bifurcation analysis
$0,1)\ni s\mapsto\left(\frac{se^{-\frac{s(1-s)}{s(1-s)-\gamma}}}{s(1-s)-\gamma},\frac{se^{-\frac{s(1-s)}{s(1-s)-\gamma}}}{1-s}\right)$
@@
The continuation algorithm terminates with a failure to converge (denoted by MX) as, near $\gamma=1/8$, $\ell_{1}$ varies rapidly between 0 at the DH point and $\infty$ at the BTP point (which are coincident in the limit of $\gamma=1/8$).
@@ 2.3. Periodic orbits
Figure 6. Bifurcation diagram for equilibria and periodic orbits with free parameters $\sigma$ and $\delta$ in a small parameter region containing the Bogdanov-Takens point $(\sigma_{\mathrm{BT,1}},\delta_{\mathrm{BT,1}})$ (subscript 11 not shown in the figure) in Fig. 3(b). For better visibility of the bifurcation curves, the coordinates of the parameter plane have been rotated counterclockwise by $\theta=150^{\circ}$ around the Bogdanov-Takens point, such that $\tilde{\sigma}=\cos\theta(\sigma-\sigma_{\mathrm{BT}})-\sin\theta(\delta-\delta_{\mathrm{BT}})$, $\tilde{\delta}=\sin\theta(\sigma-\sigma_{\mathrm{BT}})+\cos\theta(\delta-\delta_{\mathrm{BT}})$. Other parameters: $\beta=0$, $\gamma=0.1$.
For homoclinic orbits connecting to saddle-node bifurcation equilibria, the determinant (det) should equal 0, whereas the determinant is negative in the saddle case. Across the transition between such orbits, the shape of the homoclinic turns from a smooth curve to a non-smooth curve in phase space (the “corner” is the location of the saddle, see phase portraits in Fig. 5). At the transition point, the homoclinic orbit is a non-central connection to a saddle-node [2]. In the screen output, approximate detection of such a transition is accomplished by monitoring for a crossing by det of 0.05-0.05 (a suitably chosen small, negative number) and the corresponding points are denoted by NCS. Four such non-central connections to saddle-nodes are found near the Bogdanov-Takens point $(\sigma_{\mathrm{BT,1}},\delta_{\mathrm{BT,1}})$ at the end points of the segments labeled SNIC in Fig. 6. Magnification (not included here but available using the provided code) shows that the homoclinic curve is smooth across these points.
@@ 3.1. Algebraic problems
(9) $0=kv_{1}+J^{2}v_{1},\,0=v_{1}^{\mathsf{T}}v_{1}-1,\,0=\xi^{\mathsf{T}}v_{1},$
where $J=\partial_{x}f(x,p)$. (Note that the numerical implementation does not perform such elimination since if $J$ has large entries, entries in $J^{2}$ may be even larger.) Suppose that a vector $\hat{v}_{1}$ satisfies the first equation in (9). Then k-k must be an eigenvalue of $J^{2}$ with eigenvector $\hat{v}_{1}$. If $k>0$, then $\hat{v}_{1}$ cannot be an eigenvector also of $J$ (since k<0-k<0). It follows that $\hat{v}_{1}$ and $J\hat{v}_{1}$ span a two-dimensional invariant subspace of $J$ for the purely imaginary eigenvalue $\sqrt{-k}$ with corresponding complex eigenvector $\sqrt{-k}\hat{v}_{1}+J\hat{v}_{1}$. In particular, there exist locally unique scalars $\alpha$ and $\beta$ such that the vector $\alpha\hat{v}_{1}+\beta J\hat{v}_{1}$ satisfies all three of the equations in (9).
Suppose, instead, that $k<0$. Then, if $\hat{v}_{1}$ is not an eigenvector of $J$, $\hat{v}_{1}$ and $J\hat{v}_{1}$ span an invariant subspace of $J^{2}$, which is also an invariant subspace for $J$. It follows that $J$ must have the real eigenvalues $\sqrt{-k}$ and $\sqrt{-k}$, i.e., the corresponding equilibrium is a neutral saddle. There again exist locally unique scalars $\alpha$ and $\beta$ such that the vector $\alpha\hat{v}_{1}+\beta J\hat{v}_{1}$ satisfies all three of the equations in (9).
If, instead, $\hat{v}_{1}$ is an eigenvector of $J$ (by necessity corresponding to an eigenvalue $\sqrt{-k}$ or $\sqrt{-k}$), then $\hat{v}_{1}$ and $J\hat{v}_{1}$ do not span a two-dimensional invariant subspace for $J$ and there are no constraints on the other eigenvalues of $J$. Orthogonality with respect to $\xi$ here plays the role of the eigenvalue conditions in the preceding two paragraphs. In the CSTR model example, this case corresponds to the solution family (8). Such a solution family may intersect a family of neutral saddles at a unique branch point, as observed in the preceding section for the CSTR model example.
@@ 3.2. Boundary-value problems
The function $\tau\mapsto x(\tau)$ is next approximated by a continuous function $\tau\mapsto x_{\mathrm{d}}(\tau)$ with $L$ polynomial pieces on subintervals defined on a mesh $0=\tau_{0}<\tau_{1}<\ldots\tau_{L}=1$, each of degree $n_{\mathrm{deg}}$. Finally, let $\tau_{\mathrm{coll},j}$ for $j=1,\ldots,n_{\mathrm{deg}}$ be collocation parameters defining the discretization method (e.g., Gauss-Legendre nodes [12] scaled to a base interval $[0,1]$) and $P_{\mathrm{d}}$ be the projection that maps the function $\tau\mapsto y(\tau)$ onto the unique (possibly discontinuous) piecewise-polynomial function $\tau\mapsto[P_{\mathrm{d}}y](\tau)$ that is of degree $n_{\mathrm{deg}}-1$ on each mesh interval and satisfies
@@
The defining problem (12) for a saddle-node bifurcation of periodic orbits is guided by the desire to ensure that the monodromy matrix associated with the variational problem has a generalized eigenvector $y_{\mathrm{d}}(0)$ corresponding to the Floquet multiplier $1$, orthogonal to $\dot{x}_{\mathrm{d}}(0)=f(x_{\mathrm{d}}(0),p)$. In a similar fashion, other generic local bifurcations of periodic orbits, such as period-doubling bifurcations or torus bifurcations have defining problems that ensure that the monodromy matrix has an eigenvector for a desired Floquet multiplier. For period-doubling bifurcations, this is achieved by omitting the fifth component in (12) and replacing the next-to-last component with $y_{\mathrm{d}}(1)+v$, such that $y_{\mathrm{d}}(0)$ is an eigenvector corresponding to a Floquet multiplier 1-1.
@@ 3.3. Periodic orbits near homoclinics
Consider, first, the scenario shown in Fig. 5, in which a family of periodic orbits accumulate on a connecting orbit to a saddle-node bifurcation equilibrium at a saddle-node-on-invariant curve bifurcation as $\sigma\downarrow\sigma_{\mathrm{snic}}$. Here, increases in the period of nearby periodic orbits are dominated by a slowing down of the dynamics in the immediate vicinity of the ghost of the saddle-node bifurcation equilibrium as represented schematically in Fig. 9(a). This depicts a size-$\varepsilon$ neighborhood of the saddle-node bifurcation equilibrium in the center manifold (horizontal axis is $\sigma$, vertical axis is the other center manifold direction, called $x_{\mathrm{c}}$). The parabolic curve in Fig. 9(a) is the family of equilibria associated with the saddle-node bifurcation (red dashed unstable, and green solid stable). Expanding the vector field to second order about the saddle-node bifurcation, one concludes that the time for $x_{\mathrm{c}}$ to increase from $\varepsilon$ to $\varepsilon$ for $\sigma>\sigma_{\mathrm{snic}}$ varies as
@@ 4.1. A Network of Chemical Oscillators
For fixed $A$ and $\delta$, such that $11A^{2}(11-9\delta)>100\delta^{2}$, and under variations in $\epsilon$, the second pair of eigenvalues, $\kappa_{2,\pm}$, crosses the imaginary axis transversally for $\epsilon=\epsilon^{*}:=\delta/44$ at $\pm\mathrm{i}\omega$ with
@@
The equivariant Hopf bifurcation theorem permits construction of families of periodic orbits that emanate from such a bifurcation point (see [17, 18] for the general theory in the presence of symmetry, [19] for systems with general permutation symmetry, and [20] for a study of the case of four oscillators coupled in a ring topology). Periodic orbits in a system with symmetry are classified according to their spatiotemporal symmetry, here represented by a permutation composed of $K$ subcycles and an integer cycle shift for each subcycle. For example, the notation $(1\,2\,4\,3)_{4}$ applies to a periodic orbit of period $T$ that satisfies the conditions
@@
with $K=1$ and single cycle shift $4$, while $(1\,2)_{2}(3\,4)_{1}$ applies to a periodic orbit of period $T$ that satisfies the conditions
@@
with $K=2$ and cycle shifts $2$ and $1$. Let $\ell_{k}$ equal the cycle shift of the subcycle containing index $k$, such that $\ell=(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$ is the corresponding shift vector. Then, $\ell$ is $(4,4,4,4)$ for the spatiotemporal symmetry $(1\,2\,4\,3)_{4}$ in (14), while $\ell=(2,2,1,1)$ for the spatiotemporal symmetry $(1\,2)_{2}(3\,4)_{1}$ in (15). When a subcycle contains only a single index and the corresponding cycle shift equals $1$, we sometimes omit the subcycle from the representation.
@@
For a quantitative example, suppose (as in [16]) that $A=2.0$ and $B=5.9$. In this case, the equivariant Hopf bifurcation occurs for $1000\epsilon^{*}\approx 20.45$ with $\omega^{*}\approx 0.62$ and (16) is satisfied by the spatiotemporal symmetry $(1\,2\,4\,3)_{4}$. The equivariant Hopf bifurcation theorem then predicts the existence of a family of periodic orbits emanating from the bifurcation point with limiting period $2\pi/\omega^{*}\approx 10.12$ and satisfying the conditions (14). The bifurcation diagram near the equivariant Hopf bifurcation in Fig. 10 shows the corresponding branch labeled $P_{1}$ with a typical time profile in the right panel.
@@
At the equivariant Hopf bifurcation, condition (16) is also satisfied by other spatiotemporal symmetries, such as $(1\,2)_{2}(3\,4)_{1}$ and $(1)_{1}(2\,3\,4)_{1}$ (abbreviated to $(2\,3\,4)_{1}$). The equivariant Hopf bifurcation theorem consequently predicts the existence of branches of periodic orbits satisfying the conditions (15) and
@@
Figure 11. Family $P_{1}$ of periodic orbits of (13) with $(1\,2\,4\,3)_{4}$ symmetry, with further symmetry-breaking bifurcations, resulting in families $P_{4}$, $P_{5}$, $P_{6}$ with other symmetries as indicated in the legend and illustrated with time profiles.
For each of the bifurcations at $\lambda\approx 17.00$ and $\lambda\approx 2.67$, on the other hand, a single eigenvalue crosses the unit circle at $1$. In contrast to the non-symmetric, generic case, these are not saddle-node bifurcations. Indeed, as the points are also detected as branch points, denoted by BP, and under appropriate genericity assumptions, they are expected to be symmetry-breaking flip-pitchfork bifurcations, an expectation that is borne out by further computational analysis. From each emanate secondary branches of periodic orbits that are symmetric with respect to the spatiotemporal symmetry associated with the permutation $(1\,4)_{2}(2\,3)_{2}$, i.e., such that
@@
... 1 EP 1.4473e+01 5.9863e+$0$ 1.8536e+$0$
... 2 BP 1.9650e+01 7.2452e+$0$ 2.6173e+$0$
... 3 UST 1.9650e+01 7.2452e+$0$ 2.6173e+$0$
... 4 BP 2.2810e+01 8.5430e+$0$ 3.6744e+$0$
... 5 UST 2.2810e+01 8.5430e+$0$ 3.6743e+$0$
... 6 UST 2.5346e+01 1.0683e+01 4.8199e+$0$
... 7 FP 2.5346e+01 1.0683e+01 4.8199e+$0$
... 8 BP 1.5355e+01 1.2014e+01 4.0979e+$0$
... 9 UST 1.5355e+01 1.2014e+01 4.0979e+$0$
... 10 UST 1.5209e+01 1.2009e+01 4.0666e+$0$
... 11 BP 1.5209e+01 1.2009e+01 4.0666e+$0$
... 12 BP 1.4616e+01 1.1820e+01 3.9017e+$0$
... 13 UST 1.4616e+01 1.1820e+01 3.9017e+$0$
... 14 EP 1.4000e+01 1.1364e+01 3.6991e+$0$
(Recall that $\lambda=\epsilon\times 10^{3}$.) Branch switching to $P_{6}$ in Fig. 11 from the first of the branch points reduces symmetry again. Along this branch $x_{2}(t)=x_{3}(t),\,y_{2}(t)=y_{3}(t)$ but the spatiotemporal symmetry between $(x_{1},y_{1})$ and $(x_{4},y_{4})$ has been lost, as indicated by the cycle notation $(2\,3)_{1}$ (short for $(2\,3)_{1}(1)_{1}(4)_{1}$) in the legend and the typical time profiles.